OpenStudy (anonymous):

solve the following equation: a- 9/a = 8

sam (.sam.):

a-(9)/(a)=8 Least common denominator: a a^(2)-9=8a a^(2)-8a-9=0 (a+1)(a-9)=0 a+1=0_a-9=0 a=-1_a-9=0 a=-1_a-9=0 a=-1_a=9 a=-1,9

OpenStudy (anonymous):

a^2-9=8a a^2-8a-9=0 (a-9)(a+1) a= 9 or -1

OpenStudy (anonymous):

cool thanks!